OpenStudy (anonymous):

Medals awarded if all completed and method shown :) (Simultaneous Equations) 1.) 3x+2y=11 2x- y=-3 2.) 5x-7y=27 3x-4y=16 3.) 2x- y=5 \(\frac{ x }{ 4 } + \frac{ y }{ 3 }\)=2 4.) 0.4x + 3y = 2.6 x - 2y = 4.6

OpenStudy (hba):

There are a two methods to solve this : 1) Substitution method 2) Elimination method Which one would you prefer ?

OpenStudy (anonymous):

Elimination method please :)

OpenStudy (hba):

Well in the method of elimination we try to add and subract to get rid of one variable or find a particular variable. Lets see the first one, 3x+2y=11---(1) 2x- y=-3---(2) I would reccomend you to multiply eq(2) by 2 and then add those equations You will get x.

OpenStudy (anonymous):

i get x=5 but when i put it into the equation it does not work..... :(

OpenStudy (hba):

3x+2y=11 4x-2y=-6 ---------- I think you are doing it wrong :/ try doing it again :)

OpenStudy (anonymous):

cant do it :( not good at this :/

OpenStudy (hba):

You just have to add them nothing else @Chloeee_Harris

OpenStudy (anonymous):

oh ive got it now, how would i do the next question?

OpenStudy (hba):

What did you get x as ?

OpenStudy (anonymous):

5/7

OpenStudy (hba):

x=5/7 Now put x in eq(1) and find y :)

OpenStudy (anonymous):

4 3/7 :)

OpenStudy (anonymous):

how would i do the next question?

OpenStudy (anonymous):

5x-7y=27 3x-4y=16

OpenStudy (hba):

@Chloeee_Harris You got y wrong.

OpenStudy (anonymous):

it says its correct on my answer sheet,

OpenStudy (hba):

3x+2y=11 put x=5/7 3(5/7)+2y=11 @Chloeee_Harris

OpenStudy (hba):

@Chloeee_Harris 43/7 ?

OpenStudy (anonymous):

no \[4\frac{ 3 }{ 7 }\]

OpenStudy (hba):

@Chloeee_Harris Oh i got it it is 4 3/7 or 31/7 Yeah right :)

OpenStudy (anonymous):

sorry about that, thanks for helping me, could you do the second question?

OpenStudy (hba):

5x-7y=27 3x-4y=16 Now lets see this We can multiply eq(1) by 3 and eq(2) by 5 and then subract. And get y :)

OpenStudy (anonymous):

so what would be the answer to that ^

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

question 3?

OpenStudy (anonymous):

or question 4?

OpenStudy (agent0smith):

2x- y=5 \[\frac{ x }{ 4 } + \frac{ y }{ 3 }=2 \] If you want to use elimination, you could multiply the second by 8 and subtract it from the first.

OpenStudy (agent0smith):

\[8 \times ( \frac{ x }{ 4 }+\frac{ y }{ 3 })=8 \times 2\] becomes \[2x+\frac{ 8 y }{ 3 }=16\]\[2x-y=5 \] Now subtract the second from first.

OpenStudy (anonymous):